Express using a positive exponent
c^-6
(simply your answer. type a positive exponent)
: Express using a positive exponent
c^-6
(simply your answer. type a positive exponent)

Answer by Alan3354(1926)
Question 164663: Express using a positive exponent
c^-6
(simply your answer. type a positive exponent)
: Express using a positive exponent
c^-6
(simply your answer. type a positive exponent)

Answer by Mathtut(1311)

This is simply involving the negative exponent law which states that a^-x = 1/a^x and 1/a^-x = a^x therefore to express c^-6 as a positive exponent we just follow the law so c^-6 = 1/c^6 which is a expressing it as a positive exponent. The reason this works is if you take a^-x/1 and multiply both top(numerator)and bottom(denominator) by a^x on the top you get a^-x times a^x and using the multiplication law of exponents this gives us a^0 and any number raised to the zero power excluding zero itself equals 1. On the bottom 1 multiplied by a^x equals a^x. Another way to state this is when changing from a negative exponent to a positive you just invert the base and make the exponent positive.
answer: 1/c^6